OpenStudy (anonymous):

need step by step help, The graph of y = ax 2 + bx + c passes through the points (-2, 0), (0, -2), and (2, 0). Determine the solution set of 0 = ax 2 + bx + c.

OpenStudy (kaederfds):

0=4a+2b+c ..... (1) 0=4a-2b+c ....... (2) we can find that b=0 in (1) 0=4a+0+c -2=c ..... (3) (3) >> (2) a=1/2 so the solution set is a =1/2; b=0; c=(-2) ^^

OpenStudy (anonymous):

thanks(: can you halp me with this one too? \[-3/4x ^{2} +2\]

OpenStudy (anonymous):

if the equation is written in general form, what is c?

OpenStudy (kaederfds):

c=2

OpenStudy (anonymous):

okay, thank you so much:D